# Lessons Learned from a Unifying Empirical Study of Parameter-Efficient Transfer Learning (PETL) in Visual Recognition

### v2

@@ Abstract
Parameter-efficient transfer learning (PETL) has attracted significant attention lately, due to the increasing size of pre-trained models and the need to fine-tune them for superior downstream performance. This community-wide enthusiasm has sparked a plethora of approaches. Nevertheless, a systematic study to understand their performance and suitable application scenarios is lacking, leaving questions like “when to apply PETL” and “which approach to use” largely unanswered, especially in visual recognition. In this paper, we conduct a unifying empirical study of representative PETL approaches in the context of Vision Transformers (ViT). We systematically tune their hyper-parameters to fairly compare their accuracy on downstream tasks. Our study not only offers a valuable user guide but also unveils several new insights. First, if tuned carefully, different PETL approaches can obtain quite similar accuracy in the low-shot benchmark VTAB-1K. This includes simple approaches like fine-tuning the bias terms that were reported inferior. Second, though with similar accuracy, we find that PETL approaches make different mistakes and high-confidence predictions, likely due to their different inductive biases. Such an inconsistency (or complementariness) opens up the opportunity for ensemble methods, and we make preliminary attempts at this. Third, going beyond the commonly used low-shot tasks, we find that PETL is also useful in many-shot regimes — it achieves comparable and sometimes better accuracy than full fine-tuning, using much fewer learnable parameters. Last but not least, we investigate PETL’s ability to preserve a pre-trained model’s robustness to distribution shifts (e.g., a CLIP backbone). Perhaps not surprisingly, PETL approaches outperform full fine-tuning alone. However, with weight-space ensembles, the fully fine-tuned model can better balance target (i.e., downstream) distribution and distribution shift performance, suggesting a future research direction for PETL [^1] Code is available at https://github.com/OSU-MLB/PETL_Vision..
@@ 1 Introduction
Pre-training and then fine-tuning has become the standard practice to tackle visual recognition problems (Bommasani et al., 2021). The community-wide enthusiasm for open-sourcing has made it possible to access large, powerful pre-trained models learned from a gigantic amount of data, e.g., ImageNet-21K (Ridnik et al., 2021) or LAION-5B (Schuhmann et al., 2022). More research focus has thus been on how to fine-tune such large models (Yu et al., 2023a). Among existing efforts, parameter-efficient transfer learning (PETL), a.k.a parameter-efficient fine-tuning (PEFT), has attracted increasing attention lately (Han et al., 2024; Ding et al., 2023). Instead of fine-tuning the whole model (i.e., full fine-tuning) or the last fully connected layer (i.e., linear probing), PETL approaches seek to update or insert a relatively small number of parameters to the pre-trained model (Xin et al., 2024). Doing so has several noticeable advantages. First, as named, PETL is parameter-efficient. For one downstream task (e.g., recognizing bird species or car brands), it only needs to learn and store a tiny fraction of parameters on top of the pre-trained model. Second, accuracy-wise, PETL has been shown to consistently outperform linear probing and often beat full fine-tuning, as reported on the commonly used low-shot image classification benchmark VTAB-1K (Zhai et al., 2019).
@@
(b) Target distribution vs. distribution shifts
@@
Figure 1: Highlights of our insights. (a) Downstream accuracy: if tuned carefully, different PETL methods achieve similar accuracy ($\bullet$-$\bullet$ for the range from the most to the least accurate methods) and consistently outperform linear probing ($\times$) and full fine-tuning ($\blacksquare$) on VTAB-1K. (b) Distribution shift accuracy: fine-tuning a CLIP ViT-B/16, known for its generalizability across domains, with PETL on ImageNet-1K (100 samples/class) better preserves the distribution shift accuracy (Y-axis, averaged across ImageNet-V2, ImageNet-S, ImageNet-R, and ImageNet-A) than full fine-tuning, evidenced by the $\star$ points. Interestingly, weight-space ensembles (WiSE) (Wortsman et al., 2022) is applicable between PETL’s fine-tuned model and the pre-trained model ($\blacksquare$), but not as effective as applying it to the fully fine-tuned model. Details are in section 3 and section 7.
To date, a plethora of PETL approaches have been proposed, bringing in inspiring ideas and promising results. Along with this come several excellent surveys that summarize existing PETL approaches (Yu et al., 2023a; Xin et al., 2024; Ding et al., 2023). Yet, a systematic understanding of the PETL paradigm seems still missing.
@@
For example, with so many PETL approaches, there is a lack of unifying references for when and how to apply them. Though superior accuracy was reported on the low-shot benchmark VTAB-1K, there is not much discussion on how PETL approaches achieve it. Does it result from PETL’s ability to promote transferability or prevent over-fitting? The current evaluation also raises the question of whether PETL is useful beyond a low-shot scenario. Last but not least, besides superior accuracy, do existing PETL approaches offer different, ideally, complementary information?
@@
Attempting to answer these questions, we conduct a unifying empirical study of representative PETL approaches in the context of Vision Transformers (ViT) (Dosovitskiy et al., 2020). This include Low-Rank Adaptation (LoRA) (Hu et al., 2021), Visual Prompt Tuning (VPT) (Jia et al., 2022), Adapter (Houlsby et al., 2019), and ten other approaches. We systematically tune their hyper-parameters to fairly compare their accuracy on the low-shot benchmark VTAB-1K. This includes learning rate, weight decay, and approach-specific parameters like the size of PETL parameters. Besides VTAB-1K, we examine PETL approaches on full-size downstream datasets such as CIFAR-100 (Krizhevsky et al., 2009), RESISC for remote sensing image scene classification (Cheng et al., 2017), and Clevr-Distance for depth classification with synthetic data (Zhai et al., 2019; Johnson et al., 2017). We also conduct a study on ImageNet (Deng et al., 2009) and its variants with domain shifts (Hendrycks et al., 2021a; Gao et al., 2023; Hendrycks et al., 2021b; Recht et al., 2019).
We summarize our key findings and extended analyses as follows.
@@
Representative PETL approaches perform similarly on VTAB-1K, if properly implemented. This includes methods previously considered less effective, such as fine-tuning the bias terms (Zaken et al., 2022) in the pre-trained backbone and methods originally proposed for NLP, like Adapter (Houlsby et al., 2019). Among all the hyper-parameters, we find the drop path rate (Huang et al., 2016) quite important. Ignoring it (i.e., setting it to 00) significantly degrades the performance. Overall, PETL approaches consistently outperform linear probing and full fine-tuning on all 19 image classification tasks (each with $1,000$ training examples) in VTAB-1K.
@@
While similarly accurate on average, PETL approaches make different predictions. The above finding seems daunting: if existing PETL approaches all perform similarly in terms of accuracy, do we learn anything useful beyond a single approach? This is particularly worrisome given that they fine-tune the same backbone using the same downstream data. Fortunately, our analysis shows that different PETL methods learn differently from the same data, resulting in diverse prediction errors and confidence. We attribute this to their difference in inductive biases (Neyshabur et al., 2014) — they explicitly specify different parameters to be updated or inserted. This opens up the door to leverage their discrepancy for improvement, e.g., through ensemble methods (Dietterich, 2000; Zhou, 2012) or co-training (Blum & Mitchell, 1998; Balcan et al., 2004) and we provide preliminary studies.
@@
PETL is also effective in many-shot regimes. We apply PETL beyond the low-shot regime and find it effective even with ample downstream training data — PETL can be on par or outperform full fine-tuning. This suggests that varying a fraction of parameters of a properly chosen pre-trained backbone (e.g., pre-trained on ImageNet-21K (Dosovitskiy et al., 2020)) could already offer a sufficient effective capacity (Zhang et al., 2021) to reach a performant hypothesis for downstream tasks.
@@
PETL is more robust than full fine-tuning to distribution shifts, but with weight-space ensembles, the observation is overturned. We also evaluate PETL’s robustness to distribution shifts, inspired by (Wortsman et al., 2022). We consider a CLIP backbone (Radford et al., 2021), known for its superior generalizability to distribution shifts, and apply PETL to fine-tune it with ImageNet-1K. We find that PETL preserves CLIP’s generalizability (e.g., to samples from ImageNet-Sketch or ImageNet-Rendition) better than full fine-tuning. This may not be surprising. What is interesting is that the weight-space ensembles (WiSE) between the fine-tuned and pre-trained models (Wortsman et al., 2022) apply to PETL as well to further improve the robustness without sacrificing the downstream accuracy. Nevertheless, full fine-tuning with WiSE can achieve even higher accuracy in both downstream and distribution shift data than PETL, suggesting a further research direction in PETL.
@@
What lead to PETL’s success? We attempt to answer this fundamental question by analyzing the results of our study. On VTAB-1K with 19 tasks, we find two cases: 1) in some tasks, full fine-tuning outperforms linear probing, suggesting the need to update the backbone; 2) in some tasks, linear probing outperforms full fine-tuning, suggesting either the backbone is good enough or updating it risks over-fitting. The superior accuracy of PETL in both cases suggests that PETL acts as an effective regularizer during low-shot training. Still using VTAB but with ample training data, we find that for tasks in case 1), PETL is on par with full fine-tuning, suggesting that its regularization role does not prevent the fine-tuned model from learning sufficiently from the data. For tasks in case 2), PETL can surprisingly still outperform full fine-tuning, suggesting that it effectively transfers (or preserves) some useful pre-trained knowledge that full fine-tuning may wash away. In sum, PETL succeeds as a high-capacity learner equipped with an effective regularizer.
@@
Contributions. Instead of chasing the leaderboard, we systematically understand existing approaches via a unifying study. Our contribution is thus not a technical novelty, but: (1) a systematic framework enabling consistent and reproducible evaluations of PETL methods; (2) a set of empirical recommendations on when and how to use PETL methods for practitioners; (3) new insights for future research including leveraging PETL’s prediction differences and exploring robust fine-tuning.
What do we not investigate? There are many aspects that one can ask about PETL. Our study does not consider computation-specific properties like memory usage and FLOPS.
@@ 2.1 Large pre-trained models
Pre-trained models have become an indispensable part of modern AI development (Bommasani et al., 2021). Building upon neural networks with millions if not billions of parameters and gigantic amounts of training data, these large pre-trained models have led to groundbreaking results in various downstream tasks (Liang et al., 2024; Moor et al., 2023) and shown several emerging capabilities not observed previously (Khan et al., 2022; Li et al., 2024; Bommasani et al., 2021). For example, in computer vision, a Vision Transformer (ViT) (Dosovitskiy et al., 2020) trained with ImageNet-21K (around $14$M images) leads to consistent gains v.s. a ViT trained with ImageNet-1K (around $1.3$M images) (Dosovitskiy et al., 2020). ViTs pre-trained with millions of image-text pairs via a contrastive objective function (e.g., a CLIP-ViT model) (Radford et al., 2021; Cherti et al., 2023) show an unprecedented zero-shot capability and robustness to distribution shifts (Radford et al., 2021). In this paper, we focus on the ImageNet-21K-ViT and use the CLIP-ViT in a robustness study.
Vision Transformer (ViT).
We briefly review ViTs (Dosovitskiy et al., 2020), which are adapted from the Transformer-based models (Vaswani et al., 2017) in NLP. ViTs divide an image into a sequence of $N$ fixed-sized patches and treat them like NLP tokens. Each patch is first embedded into a $D$-dimensional vector ${\bm{x}}^{(n)}_{0}$ with positional encoding. The sequence of vectors is then prepended with a “CLS” vector ${\bm{x}}_{0}^{(\text{Class})}$ to generate the input $\bm{Z}_{0}=[{\bm{x}}_{0}^{(\text{Class})},{\bm{x}}^{(1)}_{0},\cdots,{\bm{x}}^{(N)}_{0}]\in\mathbb{R}^{D\times(1+N)}$ to a ViT, composed of $M$ Transformer layers. We use super-/sub-script to index token/layer. The output of the “CLS” token ${\bm{x}}_{M}^{(\text{Class})}$ is used as the image representation.
Each of the ViT’s $M$ Transformer layers consists of a multi-head self-attention (MSA) block, a multi-level perceptron (MLP) block, two Layer Normalization (LN) blocks (Ba et al., 2016), and two residual links. The $m$-th Transformer layer can be formulated as
@@
Without loss of generality, let us consider an MSA block with a single head. Given a generic input $\bm{Z}\in\mathbb{R}^{D\times(1+N)}$, this block first projects it into three matrices, Query $\bm{Q}$, Key $\bm{K}$, and Value $\bm{V}$
@@
where $\bm{W}_{Q/K/V}\in\mathbb{R}^{D\times D}$ are projection matrices. The output of this block is then formulated as
$\bm{V}\times\mathsf{Softmax}(\frac{\bm{K}^{\top}\bm{Q}}{\sqrt{D}})\quad\in\mathbb{R}^{D\times(1+N)}.$ (4)
2.2 Parameter efficient transfer learning (PETL)
@@
Fine-tuning is arguably the most common way to tailor a pre-trained model for downstream tasks. As the size of pre-trained models gets larger, copying and updating all the parameters for one downstream task becomes inefficient. PETL has thus emerged as a promising paradigm.
PETL was originally developed in NLP (He et al., 2021a; Lester et al., 2021; He et al., 2022b; Mao et al., 2022; Sung et al., 2021; Zaken et al., 2022; Asai et al., 2022; Vu et al., 2022; Liu et al., 2022a; Su et al., 2022; Zhong et al., 2022) and has attracted increasing attention in vision (Jia et al., 2022; Chen et al., 2022b; Jie & Deng, 2022; Zhang et al., 2022b; Liu et al., 2022b; Lian et al., 2022). Existing approaches can generally be categorized into four groups: prompt-based, adapter-based, direct selective parameter tuning, and efficient selective parameter tuning. We focus on visual recognition and compare representative PETL approaches applicable to ViTs. During fine-tuning, all approaches learn a new FC layer for prediction.
@@
Prompt-based approaches. Prompt-based learning emerged in NLP (Liu et al., 2023; Lialin et al., 2023). The core concept is to augment the input data with task-specific hints (prompts). Visual Prompt Tuning (VPT) (Jia et al., 2022) adapts such an idea to ViTs. Specifically, its deep version (VPT-Deep) prepends a set of soft prompts to the input tokens of each Transformer layer (i.e., $\{\bm{Z}_{m}\}_{m=0}^{M-1}$) and only optimizes the prompts during fine-tuning. Other representative works in this category include (Yu et al., 2023b; Tu et al., 2023; Gu et al., 2023).
Adapter-based approaches. This category typically introduces additional trainable parameters (e.g., an MLP block) to the frozen pre-trained model (Lialin et al., 2023). It was initially developed for multi-domain adaptation (Rebuffi et al., 2017; 2018) and continual learning (Rosenfeld & Tsotsos, 2018; Mai et al., 2022), and was subsequently extended to the NLP and vision domains to adapt Transformer-based models (Houlsby et al., 2019; Yu et al., 2023b).
@@
We consider five popular adapter-based methods. Houl. Adapter (Houlsby et al., 2019) is the first adapter-based PETL approach. It inserts two Adapters — a two-layer bottleneck-structured MLP with a residual link — into each Transformer layer, one after the MSA block and the other after the MLP block. Pfeif. Adapter (Pfeiffer et al., 2021) inserts the Adapter solely after the MLP block, a strategy shown effective in recent studies (Hu et al., 2021). AdaptFormer (Chen et al., 2022b) inserts the Adapter in parallel with the original MLP block in a Transformer layer, different from the sequential design of Houl. and Pfeif. Adapter. One can view it as an ensemble, summing the task-specific features (by the Adapter) and the task-agnostic features (by the original MLP) to form $\bm{Z}_{m}$ in Equation 2. ConvPass (Jie & Deng, 2022) introduces a convolutional-based bottleneck module (without a skip link) that explicitly encodes visual inductive biases: the 2D convolution is performed over tokens of nearby patches. The module is inserted in parallel with the MSA and/or MLP block. RepAdapter (Luo et al., 2023) introduces a linear Adapter with group-wise transformations (Luo et al., 2022) and sequentially inserts two such modules after both MSA and MLP blocks.
@@
Direct selective parameter tuning. This category selectively updates a subset of parameters of the pre-trained model, seen as a trade-off between full fine-tuning and linear probing. We consider three approaches. BitFit (Zaken et al., 2022) updates the bias terms, including those in the Q/K/V projections, the MLP blocks, the LN blocks, and the projection for patch embeddings. LayerNorm (Basu et al., 2023) updates the trainable parameters of the LN blocks in each Transformer layer. DiffFit (Xie et al., 2023) updates both the bias terms and the LN blocks and inserts learnable factors to scale the features after the MSA and the MLP blocks. Instead of updating parameters, SSF (Lian et al., 2022) linearly adapts intermediate features, motivated by feature modulation (Huang & Belongie, 2017; Perez et al., 2018). For an intermediate feature $\bm{Z}\in\mathbb{R}^{D\times(N+1)}$, SSF learns a $D$-dimensional scaling vector and a $D$-dimensional additive vector broadcasting to the tokens.
@@
Efficient selective parameter tuning. Unlike the above category which directly updates parameters, this category learns additive residuals (e.g., $\Delta\bm{W}$) to the original parameters (e.g., $\bm{W}$). By injecting a low-rank constraint to the residuals, this category effectively reduces the learnable parameters. LoRA (Hu et al., 2021), arguably the most well-known approach, parameterizes the residuals by low-rank decomposition to update the Query/Value projection matrices $\bm{W}_{Q/V}\in\mathbb{R}^{D\times D}$. Concretely, to update a $\bm{W}\in\mathbb{R}^{D\times D}$ matrix, LoRA learns $\bm{W}_{\text{down}}\in\mathbb{R}^{r\times D}$ and $\bm{W}_{\text{up}}\in\mathbb{R}^{D\times r}$ with $r\ll D$, and forms the additive residual by $\Delta\bm{W}=\bm{W}_{\text{up}}\bm{W}_{\text{down}}\in\mathbb{R}^{D\times D}$. Factor Tuning (FacT) (Jie & Deng, 2023) extends the idea of matrix decomposition into tensor decomposition. It stacks the $D\times D$ learnable matrices in all the Transformer layers into a 3D tensor and learns an additive residual parameterized by the well-established Tensor-Train (TT) (Oseledets, 2011) and Tucker (TK) (De Lathauwer et al., 2000) formulations.
More detailed descriptions of ViT and PETL methods can be found in Appendix B.
2.3 Related work and comparison
The community-wide enthusiasm for PETL has led to multiple survey articles (Yu et al., 2023a; Xin et al., 2024; Han et al., 2024). Meanwhile, several empirical, integrative, and theoretical studies were presented, mostly based on NLP tasks, attempting to provide a holistic understanding. (He et al., 2021a; Mao et al., 2021) provided unified views to methodologically connect PETL approaches. (Chen et al., 2022a; Ding et al., 2023; He et al., 2021b) and (He et al., 2022a) empirically compared PETL approaches on NLP and vision tasks, respectively, while (Fu et al., 2023) offered a theoretical stability and generalization analysis. Accuracy-wise, (Chen et al., 2022a; Ding et al., 2023; He et al., 2021b) found that PETL is robust to over-fitting and quite effective in NLP tasks under low-data regimes. This is, however, not the case for vision tasks: (He et al., 2022a) showed that representative PETL approaches like LoRA and Adapter cannot consistently outperform either full fine-tuning or linear probing. In terms of why PETL works, (Fu et al., 2023) framed PETL as sparse fine-tuning and showed that it imposes a regularization by controlling stability; (Ding et al., 2023; He et al., 2022a) framed PETL as (subspace) optimization; (Ding et al., 2023) further discussed the theoretical principle inspired by optimal control.
Our study strengthens and complements the above studies and offers new insights. First, we compared over ten PETL approaches, more than any of the above. We carefully tune the hyper-parameters, aiming to reveal the faithful accuracy of each approach. This is particularly important for the vision community because there have been no unifying references for PETL accuracy; simple approaches like BitFit have often been reported as quite inferior; the effectiveness of other approaches was reported quite discrepant from the study in NLP. Second, we go beyond a competition perspective to investigate a complementary perspective of PETL approaches. We show that different PETL approaches offer effective base learners for model ensembles. Third, we go beyond downstream accuracy to investigate PETL’s effectiveness in maintaining out-of-distribution robustness. Fourth, we systematically analyze the results from low-shot and many-shot regimes and identify two distinct patterns among PETL, full fine-tuning, and linear probing, extending the understanding of PETL.
3 How Do PETL Methods Perform in Low-Shots Regime?

### v4

# Lessons and Insights from a Unifying Study of Parameter-Efficient Fine-Tuning (PEFT) in Visual Recognition

@@ Abstract
Parameter-efficient fine-tuning (PEFT) has attracted significant attention due to the growth of pre-trained model sizes and the need to fine-tune (FT) them for superior downstream performance. Despite a surge in new PEFT methods, a systematic study to understand their performance and suitable application scenarios is lacking, leaving questions like “when to apply PEFT” and “which method to use” largely unanswered, especially in visual recognition. In this paper, we conduct a unifying empirical study of representative PEFT methods with Vision Transformers. We systematically tune their hyper-parameters to fairly compare their accuracy on downstream tasks. Our study offers a practical user guide and unveils several new insights. First, if tuned carefully, different PEFT methods achieve similar accuracy in the low-shot benchmark VTAB-1K. This includes simple approaches like FT the bias terms that were reported inferior. Second, despite similar accuracy, we find that PEFT methods make different mistakes and high-confidence predictions, likely due to their different inductive biases. Such an inconsistency (or complementarity) opens up the opportunity for ensemble methods, and we make preliminary attempts at this. Third, going beyond the commonly used low-shot tasks, we find that PEFT is also useful in many-shot regimes, achieving comparable or better accuracy than full FT while using significantly fewer parameters. Lastly, we investigate PEFT’s ability to preserve a pre-trained model’s robustness to distribution shifts (e.g., CLIP). Perhaps not surprisingly, PEFT approaches outperform full FT alone. However, with weight-space ensembles, full FT can better balance target distribution and distribution shift performance, suggesting a future research direction for robust PEFT [^1] https://github.com/OSU-MLB/ViT_PEFT_Vision.
@@ 1 Introduction
Pre-training and then fine-tuning (FT) has become the standard practice to tackle visual recognition problems [9]. The community-wide enthusiasm for open-sourcing has made it possible to access large, powerful pre-trained models learned from a gigantic amount of data, e.g., ImageNet-21K [79] or LAION-5B [81]. More research focus has thus been on how to FT such large models [101]. Among existing efforts, parameter-efficient transfer learning (PEFT), has attracted increasing attention lately [28, 17]. Instead of FT the whole model (i.e., full FT) or the last fully connected layer (i.e., linear probing), PEFT approaches seek to update or insert a relatively small number of parameters to the pre-trained model [99]. Doing so has several noticeable advantages. First, as named, PEFT is parameter-efficient. For one downstream task (e.g., recognizing bird species or car brands), it only needs to learn and store a tiny fraction of parameters on top of the pre-trained model. Second, accuracy-wise, PEFT has been shown to consistently outperform linear probing and often beat full FT, as reported on the commonly used low-shot image classification benchmark VTAB-1K [104].
@@
(b) Prediction overlaps (5K most confident)
(c) Target distribution vs. distribution shifts
@@
Figure 1: Highlights of our insights. (a) Downstream accuracy: with proper implementation and fair tuning, different PEFT methods achieve similar accuracy ($\bullet$-$\bullet$: the range from the most to the least accurate methods) and consistently outperform linear probing ($\times$) and full FT ($\blacksquare$) on VTAB-1K. (b) Diverse predictions: despite reaching similar downstream performance, different PEFT methods produce diverse predictions. This opens new opportunities for ensemble approaches and other learning paradigms (e.g. semi-supervised learning) that can exploit the prediction discrepancies. (c) Distribution shift accuracy: FT a CLIP ViT-B/16, known for its generalizability across domains, with PEFT on ImageNet-1K (100 samples/class) better preserves the distribution shift accuracy (Y-axis, averaged across ImageNet-(V2, S, R, A) than full FT, evidenced by the $\star$ points. Interestingly, weight-space ensembles (WiSE) [96] is applicable between PEFT’s FT model and the pre-trained model ($\blacksquare$), but not as effective as applying it to the fully FT model. Details are in section 3, section 4 and section 7.
@@
To date, a plethora of PEFT approaches have been proposed, bringing in inspiring ideas and promising results. Along with this come several excellent surveys that summarize existing PEFT approaches [101, 99, 17]. Yet, a systematic understanding of the PEFT paradigm seems still missing. For example, with so many PEFT approaches, there is a lack of unifying references for when and how to apply them. Though superior accuracy was reported on the low-shot benchmark VTAB-1K, there is not much discussion on how PEFT approaches achieve it. Does it result from PEFT’s ability to promote transferability or prevent over-fitting? The current evaluation also raises the question of whether PEFT is useful beyond a low-shot scenario. Last but not least, besides superior accuracy, do existing PEFT approaches offer different, ideally, complementary information?
@@
Attempting to answer these questions, we conduct a unifying empirical study of representative PEFT methods for Vision Transformers [19], including Low-Rank Adaptation (LoRA) [37], Visual Prompt Tuning (VPT) [42], Adapter [36], and ten other approaches. We systematically tune their hyper-parameters to fairly compare their accuracy on the low-shot benchmark VTAB-1K. This includes learning rate, weight decay, and method-specific parameters like the PEFT module sizes. Besides VTAB-1K, we examine PEFT methods on full-size downstream datasets such as CIFAR-100 [49], RESISC for remote sensing [12], and Clevr-Distance for depth classification [104, 45]. We also conduct a study on ImageNet [15] and its variants with domain shifts [34, 25, 35, 78] for robustness evaluation.
We summarize our key findings and analyses as follows:
@@
Representative PEFT methods perform similarly on VTAB-1K, when properly implemented and tuned (1(a)). This includes methods previously considered less effective, such as BitFit [103], which FT only the bias terms of the frozen backbone. Methods originally proposed for NLP, like Adapter [36] and LoRA [37] also exhibit impressive performance when their bottleneck dimensions are carefully tuned. Among all the hyperparameters, we find the drop path rate [38] particularly important. Ignoring it (i.e., setting it to 00) significantly degrades the performance, potentially due to overfitting. Overall, PEFT methods consistently outperform linear probing and full FT on all 19 image classification tasks (with $1,000$ training samples) in VTAB-1K.
@@
While similarly accurate on average, PEFT approaches make different predictions (1(b)). The above finding seems daunting: if existing PEFT approaches all perform similarly in terms of accuracy, do we learn anything useful beyond a single approach? This is particularly worrisome given that they FT the same backbone using the same downstream data. Fortunately, our analysis shows that different PEFT methods learn differently from the same data, resulting in diverse prediction errors and confidence. We attribute this to their inductive bias differences [70] — they explicitly specify different parameters to be updated or inserted. This opens up the door to leverage their discrepancy for improvement, e.g., through ensemble methods [16, 110] or co-training [8, 6] and we provide preliminary studies.
@@
PEFT is also effective in many-shot regimes. We extend PEFT beyond low-shot regime and find it effective even with ample downstream training data — PEFT can be on par or surpass full FT. This suggests that adjusting only a fraction of parameters in a suitably pre-trained backbone (e.g., pre-trained on ImageNet-21K [19]) could already offer a sufficient capacity [105] to reach a performant hypothesis for downstream tasks.
@@
PEFT appears more robust than full FT to distribution shifts, but WiSE overturns this advantage (1(c)). We also assess PEFT’s robustness to distribution shifts, following [96]. We consider a CLIP backbone [75], known for its superior generalizability to distribution shifts, and FT it with PEFT on ImageNet-1K. We find that PEFT retains CLIP’s generalizability (e.g., to samples from ImageNet-(V2, S, R, A)) better than full FT. This may not be surprising. What is interesting is that the weight-space ensembles (WiSE) between the FT and pre-trained models [96] is compatible with PEFT as well to further improve the robustness without sacrificing the target accuracy. To the best of our knowledge, we are the first to explore WiSE for PEFT. Nevertheless, full FT with WiSE can achieve even higher accuracy in both downstream and distribution shift data than PEFT, suggesting a further research direction in robust PEFT.
@@
What lead to PEFT’s success? We attempt to answer this fundamental question by analyzing the findings in our study. On VTAB-1K with 19 tasks, we identify two scenarios: (1) in certain tasks, full FT outperforms linear probing, suggesting the need to update the backbone; (2) in other tasks, linear probing outperforms full FT, suggesting either the backbone is good enough or updating it risks over-fitting. The superior accuracy of PEFT in both scenarios suggests that PEFT acts as an effective regularizer during low-shot training. Still using VTAB but with ample training data, we find that for scenario (1) tasks, PEFT performs similarly with full FT, suggesting that its regularization role does not impede model learning from abundant data. For scenario (2) tasks, PEFT can surprisingly still outperform full FT, indicating that it effectively transfers (or preserves) crucial pre-trained knowledge that full FT may discard. Overall, PEFT succeeds as a high-capacity learner equipped with an effective regularizer.
@@
Contributions. Instead of chasing the leaderboard, we systematically scrutinize existing methods via a unifying study. Our contribution is thus not a technical novelty, but: (1) a systematic framework for reproducible evaluations of PEFT methods; (2) a set of empirical recommendations on when and how to use PEFT methods for practitioners (section 3, section 5); (3) new insights for future research including leveraging PEFT’s prediction differences (section 4) and exploring robust fine-tuning (section 7).
@@ 2.1 Large pre-trained models
Building upon networks with millions (or billions) of parameters and massive training data, large pre-trained models have led to groundbreaking results in various downstream tasks [56, 69] and shown emerging capabilities not observed previously [47, 53, 9]. For example, a Vision Transformer (ViT) [19] trained with ImageNet-21K ($14$M images) leads to consistent gains v.s. a ViT trained with ImageNet-1K ($1.3$M images) [19]. ViTs pre-trained with millions of image-text pairs via a contrastive objective function (e.g., CLIP-ViT) [75, 13] show an unprecedented zero-shot capability and robustness to distribution shifts [75]. We focus on the ImageNet-21K-ViT and CLIP-ViT in this paper.
Vision Transformer (ViT). A ViT contains $M$ Transformer layers consisting of a multi-head self-attention (MSA) block, a multi-level perceptron (MLP) block, two Layer Normalization (LN) blocks [4], and two residual links. The $m$-th Transformer layer can be formulated as
@@
where $\bm{Z}_{m-1}$ is the output of the preceding $(m-1)$-th Transformer layer. Without loss of generality, let us consider a single-head MSA where the input $\bm{Z}$ is first projected into three matrices, $\bm{Q}$, $\bm{K}$, and $\bm{V}$. The output of this block is then formulated as:
@@
$\bm{V}\times\mathsf{Softmax}(\frac{\bm{K}^{\top}\bm{Q}}{\sqrt{D}}).$ (4)
2.2 Parameter-Efficient Fine-Tuning (PEFT)
@@
Fine-tuning is arguably the most common way to tailor a pre-trained model for downstream tasks [111, 96, 89]. As the size of pre-trained models gets larger, updating and storing all the parameters for one downstream task becomes inefficient. PEFT has thus emerged as a promising paradigm. PEFT was originally developed in NLP [29, 51, 33, 68, 84, 103, 3, 91, 58, 83, 109] and has attracted increasing attention in vision [42, 11, 43, 107, 59, 55]. Existing approaches can generally be categorized into four groups: prompt-based, adapter-based, direct selective tuning, and efficient selective tuning. We focus on visual recognition and compare representative PEFT approaches applicable to ViTs.
@@
Prompt-based. Prompt-based method emerged in NLP [57, 54] whose core concept is augmenting the input data with task-specific prompts. Visual Prompt Tuning (VPT) [42] adapts such an idea to ViTs. Its deep version (VPT-Deep) prepends a set of soft prompts to the input tokens of each Transformer layer (i.e., $\{\bm{Z}_{m}\}_{m=0}^{M-1}$) and only optimizes the prompts during FT. Other representative works in this category include [88, 102, 27, 87, 5, 94].
@@
Adapter-based. They typically introduce additional trainable parameters to a frozen pre-trained model [54]. Initially developed for domain adaptation [76, 77] and continual learning [80, 66], they have been extended to adapt Transformer-based models in NLP and vision [36, 59, 99, 108, 102]. We consider five popular methods. As the first adapter-based method, Houl. Adapter  [36] inserts two Adapters (a two-layer bottleneck-structured MLP with a residual link) into each Transformer layer, one after the MSA block and one after the MLP block. Pfeif. Adapter [74] inserts the Adapter solely after the MLP block, shown effective in recent studies [37]. AdaptFormer [11] adds an Adapter in parallel with the original MLP block, different from the sequential design of previous methods. ConvPass [43] inserts a convolutional-based Adapter that explicitly encodes visual inductive biases by performing 2D convolution over nearby patch tokens. This module is inserted parallel to the MSA and MLP blocks. RepAdapter [63] introduces linear Adapters with group-wise transformations [62], placed sequentially after MSA and MLP.
@@
Direct selective tuning. They selectively update a subset of parameters of the backbone, striking a balance between full FT and linear probing. We consider three approaches. BitFit [103] updates the bias terms, including those in the patch embeddings projection, the Q/K/V weights, the MLP and LN blocks. LayerNorm [7] updates the parameters of the LN blocks. DiffFit [97] updates both the bias terms and the LN blocks and inserts learnable factors to scale the features after the MSA and the MLP blocks. Instead of updating parameters, SSF [55] linearly adapts intermediate features, motivated by feature modulation [39].
@@
Efficient selective tuning. Instead of directly updating parameters, these methods learns additive residuals (e.g., $\Delta\bm{W}$) to the original parameters (e.g., $\bm{W}$). By injecting a low-rank constraint to the residuals, this category effectively reduces the learnable parameters. LoRA [37], arguably the most well-known approach, parameterizes the residuals by low-rank decomposition to update the weights. Concretely, to update a $\bm{W}\in\mathbb{R}^{D\times D}$ matrix, LoRA learns $\bm{W}_{\text{down}}\in\mathbb{R}^{r\times D}$ and $\bm{W}_{\text{up}}\in\mathbb{R}^{D\times r}$ with $r\ll D$, and forms the additive residual by $\Delta\bm{W}=\bm{W}_{\text{up}}\bm{W}_{\text{down}}\in\mathbb{R}^{D\times D}$. FacT [44] extends the idea of matrix decomposition into tensor decomposition. It stacks the $D\times D$ learnable matrices in all the Transformer layers into a 3D tensor and learns an additive residual parameterized by Tensor-Train (TT) [72] and Tucker (TK) [14] formulations. More detailed summary of ViT and a survey of PEFT methods can be found in Appendix B.